OpenStudy (cherrilyn):

Evaluate the integral ; integral of (x^2+x+3)dx/ (x-1)^3

OpenStudy (cherrilyn):

\[\int\limits_{}^{} (x ^{2} + x + 3) dx / (x-1)^{3}\]

OpenStudy (cherrilyn):

A/x-1 + B/(x-1)^2 + C/(x-1)^3... then

OpenStudy (anonymous):

You're doing good, you want to know how to finish?

OpenStudy (anonymous):

(x2+x+3)dx/(x−1)3 = A/x-1 + B/(x-1)^2 + C/(x-1)^3

OpenStudy (anonymous):

Multiply through by \[(x-1)^{3}\]

OpenStudy (cherrilyn):

don't I have to get rid of the fractions first ?

OpenStudy (anonymous):

Multiplying by \[(x-3)^{3}\] is a way of getting rid of the fractions.

OpenStudy (cherrilyn):

jk, thats what it does haha YEAH

OpenStudy (cherrilyn):

now find out the variables right?

OpenStudy (cherrilyn):

if x = 1, c = 5 but how do I find out the other variables

OpenStudy (anonymous):

You substitute for 5 for C into the line \[x ^{2}+x+3=A(x-1)^{2}+B(x-1)+C\]

OpenStudy (anonymous):

and expand

OpenStudy (cherrilyn):

how do you expand ?

OpenStudy (anonymous):

Multiply threw with B, expand \[(x-1)^{2}\]